Manen’s closet has three pairs of pants (black, white and green), two shirts (green and white) and two pairs of shoes (black and white).


• a) How many different outfits can be made?

• b)   Write down the sample space.


• c)   What is the probability that if you close your eyes and choose randomly, you would   

choose:

• (i) Pants and shoes with the same colour?

• (ii) Pants and shirts with the same colour?

• (iii) Pants, shoes and shirts with the same colour?

• (iv) Pants, shoes and shirts with the different colours?



1
Expert's answer
2022-05-18T09:28:13-0400

a)

"\dbinom{3}{1}\dbinom{2}{1}\dbinom{2}{1}=12"

12 different outfits can be made.


b) The sample space is


"S=\{B_1G_2B_3, B_1G_2W_3,B_1W_2B_3,B_1W_2W_3,"

"W_1G_2B_3, W_1G_2W_3,W_1W_2B_3,W_1W_2W_3,"

"G_1G_2B_3, G_1G_2W_3,G_1W_2B_3,G_1W_2W_3\}"

c)

(i)


"B_1G_2B_3,B_1W_2B_3,W_1G_2W_3,W_1W_2W_3"


"p=\dfrac{4}{12}=\dfrac{1}{3}"

(ii)


"W_1W_2B_3, W_1W_2W_3, G_1G_2B_3, G_1G_2W_3"

"p=\dfrac{4}{12}=\dfrac{1}{3}"

(iii)


"W_1W_2W_3"

"p=\dfrac{1}{12}"

(iv)


"B_1G_2W_3,W_1G_2B_3,G_1W_2B_3"

"p=\dfrac{3}{12}=\dfrac{1}{4}"